OpenStudy (anonymous):

find the equation of the derivative function f(x)=5x^4

OpenStudy (cwrw238):

use the power rule if f*x) = ax^n f'(x) = anx^(n-1)

OpenStudy (cwrw238):

here a=5and n = 4 plug these into the formula and you've got it

OpenStudy (anonymous):

so f(x)= 5(4)x^4-1

OpenStudy (anonymous):

20x^3

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

thanks!

OpenStudy (cwrw238):

yw